A magician projects balls vertically upwards with one second interval.If he throws the second ball after the first one has reached the max.height,the max.height is proportional to

Grade:11

1 Answers

Chetan Mandayam Nayakar
13 years ago

 

Dear Chakrala, 

let v be the speed with which balls are projected upwards. v=gt, t=1sec, therefore v=9.8m/s, max height= v2/2g=9.8*9.8/(2*9.8)=4.9m

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar respectively : Click here to download the toolbar.. 

Askiitians Expert

Chetan Mandayam Nayakar – IIT Madras